Find the minimum distance from the point (4, 2) to the parabola y²=8x.

1
Expert's answer
2021-11-03T06:49:10-0400

Solution:

Given, "y^2=8x" and point is (4, 2).

Let y=a and put in above equation.

"a^2=8x
\\\Rightarrow x=\dfrac{a^2}8"

so, the point "(\dfrac{a^2}8,a)" lies on the curve.

Let the distance between "(\dfrac{a^2}8,a)" and (4, 2) be given by D.

Then,

"D^2=(\dfrac{a^2}8-4)^2+(a-2)^2
\\ \Rightarrow D^2=\dfrac1{64}(a^2-32)^2+(a-2)^2\ ...(i)"

Differentiating both sides of (i) w.r.t "a",

"2D.D'=\dfrac1{32}(a^2-32)(2a)+2(a-2)"

We need minimum distance, so put "D'=0"

"0=\dfrac1{32}(a^2-32)(2a)+2(a-2)
\\ \Rightarrow 0=\dfrac1{16}a(a^2-32)+2(a-2)
\\ \Rightarrow 0=a(a^2-32)+32(a-2)
\\ \Rightarrow 0=a^3-32a+32a-64
\\ \Rightarrow 0=a^3-64
\\ \Rightarrow a=4"

Put this value in (i).

"D^2=\dfrac1{64}(4^2-32)^2+(4-2)^2
\\ \Rightarrow D^2=\dfrac1{64}(-16)^2+(2)^2
\\ \Rightarrow D^2=\dfrac1{64}(256)+4
\\ \Rightarrow D^2=4+4=8
\\ \Rightarrow D=2\sqrt2\ units"